A company reported the following:

Net income 270,000

Preferred Dividends 10,000

Shares of common stock outstanding 20,000

Market price per share of common stock 36.40

Calculate the company's price-earnings ratio. Round your answer to one decimal place.
Answer #1

Price Earning ratio = Market price per Share÷ Earnings per share
Earnings per Share
= (Net Income - Preferred Dividend )÷ No. of Equity Shares
($270,000 -$10,000)÷20,000 =$260,000÷20000=13

=36.40÷13

=2.8